OpenStudy (anonymous):

What is the 20th term of the sequence that begins -2, 4, -8, 16...?

OpenStudy (campbell_st):

well what is the common ratio r...?

OpenStudy (texaschic101):

first you need to figure out if this is a geometric sequence or an arithmetic sequence. Can you do that ?

OpenStudy (anonymous):

No, I don't know how

OpenStudy (anonymous):

I don't know campbell_st

OpenStudy (campbell_st):

ok... try looking at the terms -2 x ? = 4 4 x ? = -8 -8 x ? = 16 what do you think the value of ? is

OpenStudy (anonymous):

-2

OpenStudy (texaschic101):

an arithmetic sequence....the numbers are determines by adding a geometric sequence is determined by multiplying what you have is a geometric sequence and you need to find the common ratio by dividing the second term by the first term. 4 / -2 = ? correct...the ratio is -2 between each number

OpenStudy (anonymous):

Okay...

OpenStudy (campbell_st):

great, thats the common ratio r... the formula for a term in an geometric series is \[a_{n} = a \times r^{n - 1}\] in your question n = 20 and a = -2 as well as r = -2 so you get \[a_{20} = -2 \times (-2)^{20-1}\] which simplifies to \[a_{20} = -2 \times (-2)^{19}\] so put the right hand side of the equation into a calculator... as its written to find the 20th term

OpenStudy (texaschic101):

now we will use the formula: an = a1 * r^(n-1) n = the term you want to find = 20 a1 = first term = -2 r = common ratio = -2 campbell_st.....you type fast...lol

OpenStudy (anonymous):

Wow thanks, that makes perfect sense Campbell_st! I got 1,048,576

OpenStudy (anonymous):

Thank you so much texaschic101!

OpenStudy (texaschic101):

no problem...I didn't really do that much.. Read this...it will help http://www.regentsprep.org/Regents/math/algtrig/ATP2/GeoSeq.htm

OpenStudy (anonymous):

Oh you did :) Ok I will, thanks

OpenStudy (texaschic101):

:)